Suppose 100ml of a 0.2 M ZnCl2 solution is mixed with 300ml of 0.1 KNO3. What is the pH?


1
Expert's answer
2022-02-19T06:53:15-0500

ZnCl2 + KNO3

ZnCl2 = (100×0.2) = 20m mol

KNO3 = (300×0.1) = 30m mol

Total Volume= (100+300) = 400ml

Mk+ = (30÷400) = 0.075M

MNO3- = (30÷400)= 0.075M

MZn+ = (20÷400)= 0.05M